OpenStudy (dtan5457):

What;s the min and max of y=tan2x-3?

OpenStudy (zarkon):

doesn't have one

OpenStudy (zarkon):

unless that 2 is a square

OpenStudy (dtan5457):

tan doesn't have min and max? sin and cos do, through right?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):



OpenStudy (dtan5457):

i know that, just not sure if theres a min or max like cos and sin with the amplitude

ganeshie8 (ganeshie8):

I think we can say the amplitude is not finite.

OpenStudy (dtan5457):

but it just repeats itself so that would technically be a max point

OpenStudy (dtan5457):

but i can see why tangent never ends

ganeshie8 (ganeshie8):

yes \(\tan x\) is periodic with a period of \(\pi\)

ganeshie8 (ganeshie8):

\[\tan(x)=\tan(x+\pi)\]

OpenStudy (dtan5457):

so basically i can put any points i want for the y axis

ganeshie8 (ganeshie8):

Yes the function takes all real numbers as y value. Its range is \((-\infty, ~\infty)\)